# Trajectory Prediction for Autonomous Driving: Progress, Limitations, and Future Directions

### v1

@@ 4.1.1 Map-related knowledge
One spotlight example is VectorNet, proposed by Gao et al. [410], which represents HD map elements and agent trajectories in a vectorized format. Each map element is broken into polylines composed of individual vectors, forming local graphs where nodes represent vectors and edges capture spatial or sequential relationships. These local graphs are aggregated into a global graph to encode lane-to-lane or agent-to-lane interactions using a hierarchical GNN. This approach enables interaction-aware feature encoding for downstream tasks. Next, Liang et al. [377] introduced LaneGCN, which leverages HD maps by constructing a lane graph where each lane segment is a node, and edges denote connections like predecessors, successors, and neighboring lanes. The model dynamically encodes the scene by capturing the topology and dependencies of the lane graph. By integrating actor-map interactions through a fusion network, it combines real-time traffic information with map-based context for prediction. Deo et al. [483] proposed a complementary approach, which focuses on conditioning trajectory predictions on specific paths within the lane graph. The proposed method uses discrete policy rollouts to select plausible routes, reducing computational complexity by targeting relevant map regions. By decoupling lateral variability (e.g., route choices) and longitudinal variability (e.g., speed, acceleration) and focusing predictions on sampled paths, the model enhances multimodal diversity, scene compliance, and trajectory prediction.
@@ 4.1.3 Traffic-related knowledge
Another type of traffic-related knowledge is interaction knowledge, which helps participants coexist in dynamic environments. In the study [485], pairwise interactions at intersections are categorized into discrete types such as ignoring, going, or yielding, based on the behavior of the target vehicle. Ban et al. [488] extend this approach by introducing follow and yield interactions, requiring models to explicitly predict these labels to better capture traffic dynamics. For more complex scenarios, such as roundabouts, Hu et al. [486] define interaction types using reference paths of interacting vehicles, enabling the generation of conflict-free trajectories. Ju et al. [487] link interactions to acceleration patterns, proposing a sub-network that predicts vehicle acceleration to infer interaction-driven trajectories. Additionally, social norms and informal yielding practices can also be of use, e.g., pedestrians giving way to those with trolleys, or a driver signaling a pedestrian to pass. These implicit forms of interaction knowledge can be formalized and incorporated into models to better predict complex and multimodal interactions in dense traffic scenarios.
@@ 5 Accounting for Uncertainty
Figure 11: Types of Uncertainty: the left side shows the categorization followed in the section and the right part demonstrates the flow of the prediction model and propagation of perception and tracking modules’ errors.
Motion prediction in complex traffic scenarios is fraught with high degrees of uncertainty coming from both external environment and internal system. Uncertain intentions of traffic participants, randomness of human behavior, and violations of traffic rules can be considered external sources of uncertainty. Non-obedience to traffic rules can tangibly disrupt the predictability of motion, as traffic rules are meant to constrain behavior and reduce the set of possible trajectories. As for internal systems, uncertainty arises primarily from the limitations of the prediction model itself. These limitations include biases in the training data, lack of robustness to unseen scenarios, and an inability to generalize effectively across diverse driving conditions.
For the detect-track-predict paradigm, there is a notion of propagated uncertainty. Failures in the perception module, such as inaccurate bounding boxes, missed detections, or false detections, can directly degrade tracking accuracy. Errors in the tracking system, such as incorrect object associations or ghost tracklets, propagate to the prediction module, which relies on these tracklets to generate future trajectory forecasts. These cascading errors compound the uncertainty in motion forecasting. The majority of existing prediction models assume error-free detectors and trackers, typically taking ground-truth past trajectories as their input during development and evaluation. However, in real-world deployment, such assumptions are unrealistic. Each subsequent system in the pipeline should incorporate mechanisms to mitigate the effects of upstream errors and quantify uncertainties. This includes developing models that explicitly account for propagated uncertainty, ensuring the prediction module remains robust to imperfections in the detection and tracking stages.
@@
To address these challenges, uncertainty can be categorized into aleatoric and epistemic. Aleatoric uncertainty arises from the intrinsic randomness in data. In the context of autonomous driving, it is often linked to sensors’ characteristics, such as noise in sensor measurements, limited range, calibration issues, or low resolution. These factors introduce noise and ambiguities that pose inherent challenges to prediction models. Aleatoric uncertainty reflects the limitations in fully capturing the environment. For instance, dynamic scenarios, such as predicting whether a pedestrian will cross a crosswalk or continue along a sidewalk, inherently involve aleatoric uncertainty due to their multimodal nature. This uncertainty can be further categorized as heteroscedastic (varying across samples, e.g., lighting conditions affecting image quality) or homoscedastic (uniform across samples, e.g., consistently miscalibrated sensors). Epistemic uncertainty arises from the lack of comprehensive data representing the diverse scenarios that occur in real-world traffic. It reflects the limitations in the model’s knowledge and can be reduced through improved data collection, better modelling techniques, and increased training diversity. For example, cross-dataset domain adaptation is critical for addressing epistemic uncertainty. Many motion prediction models assume that training, testing, and deployment data are identically distributed, but this assumption rarely holds in real-world applications. This domain gap can lead to degraded performance in operational settings, particularly when encountering rare or unseen scenarios. Addressing epistemic uncertainty is vital to ensuring that models generalize well and perform reliably in diverse, real-world settings.
@@
In general, accounting for different sources of uncertainty in prediction models and autonomous driving, as well as quantifying and mitigating the associated risks, is crucial for the correct operation of autonomous vehicles in unfamiliar environments. This is especially important in the presence of abnormal behavior, sensor failures, partial occlusions, and complex traffic scenarios. Figure 11 summarizes the discussed categorization of uncertainty based on its sources, nature, and the phenomenon of cascaded uncertainty. For a more detailed examination of uncertainty sources and mitigation approaches in autonomous driving, we refer the reader to the review [495].
@@ 6 Planning-conditioned Trajectory Prediction
Planning-conditioned trajectory prediction methods rely on the assumption that agents act rationally and that their future trajectories are guided by a motion planner. Recent SOTA models [514, 515, 516, 517] effectively adopt this paradigm by introducing planning-prediction-coupled frameworks, where predictions are conditioned on the observed agents’ motion plans or intentions. These methods go beyond solely relying on past motion, contextual, and interaction features; by modelling the notion that agents follow potentially planned trajectories or have specific destinations to achieve. This approach significantly enhances prediction accuracy, particularly in interactive and complex scenarios.
@@ 6.1 Planning Methods for Trajectory Prediction
Motion planning literature has extensively developed methods for generating safe and efficient trajectories for ego vehicles. This research has inspired the application of motion planning algorithms to predict the trajectories of other detected agents, under the assumption that their future paths align with specific planning objectives. As a result, planning-conditioned trajectory prediction approaches adopt a categorization similar to that of motion planning literature, namely, model-based and learning-based methods [518]. Model-based approaches rely on explicit rules, physical constraints, and optimization techniques to simulate rational agent behavior. These methods are interpretable and often incorporate principles like kinematics or dynamics to ensure physically feasible predictions. On the other hand, learning-based approaches infer plans or intentions directly from observed data, enabling adaptability to complex and diverse scenarios. Combined methods integrate the strengths of both paradigms. Reviewed works of these approaches in trajectory prediction are summarized in Table 16.
@@
Type Reference Description Model-based Gong et al. [519] (2011) Optimizes a handcrafted cost function to plan paths for the observed agents. Yan et al. [520] (2014) Introduces a probabilistic motion model towards specific goals in the environment. Vasishta et al. [289] (2017) Uses A* algorithm with the potential cost function for trajectory prediction. Cui et al. [491] (2020) Accounts for planning kinematic constraints while predicting trajectories. Learning-based Previtali et al. [104] (2016) Uses IRL for pedestrian trajectories by modelling their interactions with the environment, such as avoiding collisions or heading toward crosswalks. Choi et al. [478] (2020) GAIL for matching long-term distributions over states and actions. Zhou et al. [182] (2023) Uses reusable scene encoding and a query-based decoder with adaptive refinement to handle multimodal future behaviours effectively. Chen et al. [521] (2024) Uses GNN for trajectory predictions conditioned on guided-goals. Combined Approaches Song et al. [514] (2021) (PRIME) a prediction framework combining model-based trajectory generation with learning-based evaluation to ensure accurate, feasibility-guaranteed multimodal outputs. Chen et al. [522] (2024) Combines model predictive control with learned prediction models to optimize joint behavior.
@@ 6.1.1 Model-based planning approaches
Model-based planning for trajectory prediction forecasts agents’ future paths by leveraging models that simulate dynamics, kinematic constraints, decision-making processes, and interactions with the environment. This approach integrates physical motion principles, behavioral insights, control objectives, and road structure information to predict trajectories in a structured and interpretable manner. By explicitly modelling system dynamics, these methods produce feasible trajectories that reflect agents’ capabilities and account for their interactions with the environment [523, 524]. The main components of model-based planning for trajectory prediction include dynamics models, which simulate the evolution of an agent’s state over time using kinematic or physics-based representations [525]; cost function defining optimization objective, such as minimizing energy consumption, travel time, or collision risks [519]; planner solving the optimization problem and generating trajectories that minimize the cost function while satisfying constraints.
@@ 6.1.2 Learning-based planning approaches
Learning-based planning for trajectory prediction employs data-driven models to generate efficient and adaptable trajectories for observed dynamic agents [514]. These methods utilize techniques such as reinforcement learning and imitation learning to infer planning strategies directly from data, enabling agents to navigate complex and interactive environments effectively. Training on large datasets allows learning-based approaches to capture both high-level decision-making processes (e.g., navigating traffic junctions or performing lane changes) and low-level control actions (e.g., reacting to obstacles). Unlike traditional model-based methods that rely on explicit motion models, learning-based planning adapts to diverse and unstructured scenarios by learning implicit patterns and behaviors from data. Additionally, these methods excel in trajectory prediction by modelling interactions among agents within dynamic environments, offering robust and versatile predictions [61, 411, 64].
Chen et al. [522] proposed the Interactive Joint Planning (IJP) framework, which optimizes the behaviors of the ego vehicle and surrounding agents through deep-learning-based prediction models trained on agent interaction data. This bidirectional integration ensures that the ego vehicle’s planning influences the prediction of other agents’ behaviors while incorporating the impact of their anticipated responses. The IJP framework employs Model Predictive Control (MPC) to generate precise motion plans by integrating deep-learned prediction models with gradient-based trajectory optimization techniques, enabling the vehicle to adjust its trajectory based on predicted interactions with surrounding agents. Previtali et al. [104] employ IRL to model agent behavior by inferring cost functions from observed trajectories. IRL is used to predict pedestrian trajectories by modelling their interactions with the environment, such as avoiding collisions or heading toward crosswalks. The approach involves learning a cost function that represents human decision-making processes and using a planning algorithm to generate plausible future paths based on these inferred behaviors.
The work by Deo and Trivedi [251] assumes that each surrounding agent can perform one of several possible maneuvers, such as lane changes or braking, which are learned using LSTMs without explicitly identifying their final goals. The model assigns probabilities to these potential maneuvers and predicts the corresponding trajectories by considering the interactions between the ego vehicle and surrounding agents. Next, Ziebart et al. [526] proposed a model that assumes pedestrians behave as rational planners, efficiently moving toward their destinations. The prediction model learns a cost function that explains observed pedestrian trajectories and uses it to predict future positions. The ego vehicle’s planning module integrates this predictive model to generate a trajectory that avoids pedestrian obstacles while pursuing its own objectives. The intention-interaction graph method proposed by Li et al. [527] jointly models pedestrians’ intentions (goals) and their social interactions. The graph represents each pedestrian and their intended destination as nodes, enabling the model to predict trajectories by inferring goals based on observed behaviors and interactions, without requiring explicit knowledge of the final destinations.
@@ 6.2 Anchor-based Trajectory Prediction
Goal anchor-based methods focus on predicting a set of feasible goal points that serve as anchors for trajectory generation [528]. These methods predict potential goals based on the current context and agent behavior [528, 529]. Once the goals are identified, trajectories are generated by conditioning on these anchors, often through a path-planning or motion-generation module. Examples include predicting key destinations, such as parking spots or intersections, and using goal-conditioned policies to create paths leading to these predicted goals. A significant advantage of goal anchor-based methods is their natural alignment with goal-driven tasks, such as reaching specific destinations. Trajectory anchor-based methods, on the other hand, rely on predefined or learned trajectory templates, referred to as anchors, which represent potential future paths [3]. These methods involve selecting or predicting one or more trajectory anchors as the basis for possible future motions. The anchors provide a coarse representation of the motion, which can either be refined or directly used for trajectory prediction. Examples of trajectory anchors include predefined templates for straight, curved, or turning paths. Table 17 provides a summary of reviewed anchor-based prediction methods.
@@
Type Reference Description Goal Anchor Based Zhang et al. [530] (2020) Selects potential goal points based on observed trajectories and contextual features and ranks candidate goals based on learned scoring function. Tran et al. [531] (2020) Dual-channel framework: a goal channel predicts potential destinations, and a trajectory channel generates detailed paths conditioned on these goals. Mangalam et al. [532] (2020) Generates goals by sampling plausible future endpoints directly from the scene context, leveraging map features, and observed trajectories. Mangalam et al. [533] (2021) Hierarchically models goals, waypoints, and paths to address epistemic and aleatoric uncertainty, generating scene-compliant multimodal predictions. Zhao and Wildes [534] (2021) Leverages Dynamic Time Warping (DTW) to retrieve pseudo goal positions from an expert repository of previously observed trajectories. Yao et al. [179] (2021) Uses a recognition network to encode the relationship between past trajectories and target goals, and a prior network for sampling latent variables. Wang et al. [535] (2022) Generates coarse stepwise goals iteratively, incorporates them into the encoder and decoder using an attention-based aggregator. Chiara et al. [285] (2022) U-Net-based architecture to predict future goals as probability distributions based on scene semantics. Wang et al. [536] (2023) Uses a ”GoICrop” module to aggregate map features around potential goal areas, incorporating spatial interactions between actors and maps. Dong et al. [537] (2023) The Generates goals by learning sparse representative instances of future trajectories using a mask generated from observed trajectories. Bae and Jeon [538] (2023) Predicts a sequence of control points to divide a pedestrian’s trajectory into three sections, capturing intermediate decisions. Wang et al. [528] (2023) Generates goal anchors via self-attention mechanism on the unified agent-centric scene representation, embedding rich contextual information. Chen et al. [521] (2024) Goals are predicted for each agent using a goal-guided state refinement (SSR) module. Trajectory Anchor Based Song et al. [514] (2021) Samples dynamic reference paths and imposes explicit kinematic and environmental constraints to produce feasibility-guaranteed trajectories. Varadarajan et al. [3] (2022) Learns latent anchor embeddings to guide context-aware trajectory distributions, enhancing diversity and prediction accuracy by dynamically adapting to the scene context. Zhou et al. [529] (2023) Uses a proposal module to generate adaptive trajectory anchors recurrently, which are subsequently refined by an anchor-based module, incorporating context-specific biases.

### v2

@@ 3.5.2 Imitation Learning
3.5.3 Strengths and Limitations of RL Methods
RL-based methods learn directly from expert driving demonstrations, enabling the model to replicate realistic motion patterns observed in human behavior. Imitation Learning and generative RL frameworks via stochastic policies or latent variable models naturally achieve multimodal trajectory prediction. Additionally, POMDP-based extensions allow forecasting under partial observability, where agents must reason about latent intent or unobserved factors. As for limitations, RL methods typically require extensive expert data and are computationally demanding to train. Although in IRL, the reward function is learned from demonstrations, learning a function which captures not just safety constraints but also social dynamics and traffic agents intent ambiguity is challenging to formalize. Additionally, the learned reward and policy representations can be hard to interpret when debugging is needed.
@@ 4.1.1 Map-related knowledge
One spotlight example is VectorNet, proposed by Gao et al. [407], which represents HD map elements and agent trajectories in a vectorized format. Each map element is broken into polylines composed of individual vectors, forming local graphs where nodes represent vectors and edges capture spatial or sequential relationships. These local graphs are aggregated into a global graph to encode lane-to-lane or agent-to-lane interactions using a hierarchical GNN. This approach enables interaction-aware feature encoding for downstream tasks. Next, Liang et al. [374] introduced LaneGCN, which leverages HD maps by constructing a lane graph where each lane segment is a node, and edges denote connections like predecessors, successors, and neighboring lanes. The model dynamically encodes the scene by capturing the topology and dependencies of the lane graph. By integrating actor-map interactions through a fusion network, it combines real-time traffic information with map-based context for prediction. Deo et al. [480] proposed a complementary approach, which focuses on conditioning trajectory predictions on specific paths within the lane graph. The proposed method uses discrete policy rollouts to select plausible routes, reducing computational complexity by targeting relevant map regions. By decoupling lateral variability (e.g., route choices) and longitudinal variability (e.g., speed, acceleration). Focusing predictions on sampled paths, the model enhances multimodal diversity, scene compliance, and trajectory prediction.
@@ 4.1.3 Traffic-related knowledge
Another type of traffic-related knowledge is interaction knowledge, which helps participants coexist in dynamic environments. In the study [482], pairwise interactions at intersections are categorized into discrete types such as ignoring, going, or yielding, based on the behavior of the target vehicle. Ban et al. [485] extend this approach by introducing follow and yield interactions, requiring models to explicitly predict these labels to better capture traffic dynamics. For more complex scenarios, such as roundabouts, Hu et al. [483] define interaction types using reference paths of interacting vehicles, enabling the generation of conflict-free trajectories. Ju et al. [484] link interactions to acceleration patterns, proposing a sub-network that predicts vehicle acceleration to infer interaction-driven trajectories. Additionally, social norms and informal yielding practices can also be of use, e.g., a driver signaling a pedestrian to pass. These implicit forms of interaction knowledge can be formalized and incorporated into models to better predict complex and multimodal interactions in dense traffic scenarios.
@@ 5 Accounting for Uncertainty
Figure 11: Types of Uncertainty: the left side shows the categorization followed in the section and the right part demonstrates the flow of the prediction model and propagation of perception and tracking modules’ errors in detect-track-predict paradigm.
External uncertainty arises from the environment and includes factors beyond the model’s control. These encompass unobservable agent intentions, occlusions that limit visibility, environmental variability such as adverse lighting or weather conditions, and unexpected objects or rare events, including road accidents. For example, violations of traffic rules, such as crossing solid lanes or running red lights, can severely impact prediction performance by expanding the set of plausible future trajectories and degrading model reliability. Another challenge arises from rare events, like animals crossing the road are often underrepresented in training data, leading to potential failures in scenarios that deviate from typical urban driving conditions.
Internal uncertainty stems from limitations within the prediction system itself. It reflects challenges such as limited generalization to unseen scenarios, biases in the training data, and restricted model capacity to fully capture the complexity of real-world driving behavior. This form of uncertainty exists even when the inputs are ideal, due to incomplete knowledge or overly rigid modeling assumptions. In modular pipelines, one common source of internal uncertainty is propagated error — when imperfections in earlier modules affect downstream predictions. For instance, in the detect–track–predict setup, inaccurate detections or tracking failures can distort the input to the prediction module. However, most models are trained on clean, ground-truth trajectories and assume ideal upstream performance. In practice, this assumption breaks down, and prediction models must be designed to handle noisy inputs or account for upstream errors explicitly. More generally, any system that relies on outputs from other preceding modules is vulnerable to accumulated uncertainty along the pipeline.
@@
By nature, uncertainty can be categorized into aleatoric and epistemic types. Aleatoric uncertainty arises from intrinsic randomness in the data. In the context of autonomous driving, it is often linked to sensor characteristics, such as measurement noise, limited sensing range, calibration errors, or low resolution. These factors introduce ambiguities that pose inherent challenges to prediction models. Aleatoric uncertainty reflects the inherent difficulty in fully capturing the environment. For instance, predicting whether a pedestrian will cross at a crosswalk or continue walking along the sidewalk involves uncertainty due to the multimodal nature of human behavior. Aleatoric uncertainty can be further categorized as heteroscedastic (varying across samples, e.g., lighting conditions affecting image quality) or homoscedastic (uniform across samples, e.g., consistently miscalibrated sensors). Epistemic uncertainty, on the other hand, arises from limited knowledge of the environment, often due to insufficient or non-diverse training data. It can be reduced through improved data collection, better modeling techniques, and increased exposure to diverse scenarios. For example, cross-dataset domain adaptation is critical for addressing epistemic uncertainty. Many motion prediction models assume that training, testing, and deployment data share the same distribution, but this assumption rarely holds in real-world applications. The resulting domain gap can degrade performance, especially when encountering rare or previously unseen scenarios. Addressing epistemic uncertainty is therefore essential for improving model generalization and ensuring reliable operation in diverse, real-world settings.
@@
Accounting for different sources of uncertainty and designing uncertainty-aware prediction models is an active area of research. Quantifying and mitigating uncertainty is essential for reducing associated risks and ensuring the safe operation of autonomous vehicles. Figure 11 summarizes the categorization of uncertainty based on its sources and nature. This section further discusses techniques for both mitigating and quantifying uncertainty. For a more detailed examination of uncertainty sources and mitigation approaches in autonomous driving, we refer the reader to the review by Yang et al. [492].
@@ 6 Planning-conditioned Trajectory Prediction
Planning-conditioned trajectory prediction methods rely on the assumption that agents act rationally and that their future trajectories are guided by a motion planner. Recent SOTA models [511, 512, 513, 514] effectively adopt this paradigm by introducing planning-prediction-coupled frameworks, where predictions are conditioned on the observed agents’ motion plans or intentions. These methods go beyond solely relying on past motion, contextual, and interaction features; by modelling the notion that agents follow potentially planned trajectories or have specific destinations to achieve.
@@ 6.1 Planning Methods for Trajectory Prediction
Motion planning literature has extensively developed methods for generating safe and efficient trajectories for ego vehicles. This has inspired the application of planning algorithms to predict the trajectories of other agents, under the assumption that their future paths align with specific objectives. Accordingly, planning-conditioned trajectory prediction methods adopt a similar categorization to motion planning: model-based, learning-based, and combined approaches [515]. Model-based approaches rely on explicit rules, physical constraints, and optimization techniques, often using kinematic or dynamic models to ensure physical plausibility. In contrast, learning-based approaches infer plans or intentions directly from observed data. Combined methods aim to integrate the strengths of both paradigms. Reviewed works using these approaches for trajectory prediction are summarized in Table 16.
@@
Type Reference Description Model-based Gong et al. [516] (2011) Optimizes a handcrafted cost function to plan paths for the observed agents. Yan et al. [517] (2014) Introduces a probabilistic motion model towards specific goals in the environment. Vasishta et al. [286] (2017) Uses A* algorithm with the potential cost function for trajectory prediction. Cui et al. [488] (2020) Accounts for planning kinematic constraints while predicting trajectories. Learning-based Previtali et al. [104] (2016) Uses IRL for pedestrian trajectories by modelling their interactions with the environment, such as avoiding collisions or heading toward crosswalks. Choi et al. [475] (2020) GAIL for matching long-term distributions over states and actions. Zhou et al. [179] (2023) Uses reusable scene encoding and a query-based decoder with adaptive refinement to handle multimodal future behaviour. Chen et al. [518] (2024) Uses GNN for trajectory predictions conditioned on guided-goals. Combined Approaches Song et al. [511] (2021) PRIME combines model-based generation with learning-based evaluation to ensure accuracy, feasibility, and multimodality in trajectory prediction. Chen et al. [519] (2024) Combines model predictive control with learned prediction model to optimize joint behavior.
@@ 6.1.1 Model-based planning approaches
Model-based planning for trajectory prediction forecasts agents’ future paths by leveraging models that simulate dynamics, kinematic constraints, decision-making processes, and interactions with the environment. These approaches integrate physical motion principles, behavioral models, control objectives, and road structure information to generate interpretable and feasible trajectories that reflect agents’ capabilities [520, 521]. Core components include: dynamics models representing how agents’ state evolves over time based on kinematic equations or physical motion laws [522]; cost functions, which define optimization objectives such as minimizing energy use, travel time, or collision risks [516]; and planners used to solve the resulting optimization problems under physical and contextual constraints.
@@ 6.1.2 Learning-based planning approaches
Learning-based planning for trajectory prediction employs data-driven models to generate trajectories for observed dynamic agents [511]. These methods often use reinforcement learning to infer planning strategies directly from data. By training on large datasets, they capture both high-level decision-making (e.g., navigating junctions or changing lanes) and low-level control behaviors (e.g., reacting to obstacles). Unlike traditional model-based methods that depend on explicit motion models, learning-based planning adapts to diverse, unstructured environments by learning implicit behavioral patterns. Moreover, by modeling multi-agent interactions, these approaches produce robust predictions in dynamic and complex traffic scenarios [61, 408, 64].
Previtali et al. [104] employ IRL to model agent behavior by inferring cost functions from observed trajectories. IRL is used to predict pedestrian trajectories by modelling their interactions with the environment, such as avoiding collisions or heading toward crosswalks. The approach involves learning a cost function that represents human decision-making processes and using a planning algorithm to generate plausible future paths based on these inferred behaviors. Further, the study by Deo and Trivedi [248] assumes that each surrounding agent can perform one of several maneuvers, such as lane changes or braking, which are learned using LSTMs, and predicts the corresponding trajectories by assigning probabilities to these maneuver classes.
Next, Ziebart et al. [523] proposed a model that assumes pedestrians behave as rational planners, efficiently moving toward their destinations. The prediction model learns a cost function that explains observed pedestrian trajectories and uses it to predict future positions. The ego vehicle’s planning module integrates this predictive model to generate a trajectory that avoids pedestrian obstacles while pursuing its own objectives. The intention-interaction graph method proposed by Li et al. [524] jointly models pedestrians’ intentions (goals) and their social interactions. The graph represents each pedestrian and their intended destination as nodes, enabling the model to predict trajectories by inferring goals based on observed behaviors and interactions, without requiring explicit knowledge of the final destinations.
@@ 6.2 Anchor-based Trajectory Prediction
Goal anchor-based methods focus on predicting a set of feasible goal points that serve as anchors for trajectory generation [525]. These methods predict potential goals based on the current context and agent behavior [525, 526]. Once the goals are identified, trajectories are generated by conditioning on these anchors, often through a path-planning or motion-generation module. Examples include predicting key destinations, such as parking spots or intersections, and using goal-conditioned policies to create paths leading to these predicted goals. A significant advantage of goal anchor-based methods is their natural alignment with goal-driven tasks, such as reaching specific destinations. Trajectory anchor-based methods rely on predefined or learned trajectory templates that represent potential future paths [3]. These templates serve as coarse motion representations and are selected or predicted as bases for future trajectories. The anchors can either be directly used or refined further during prediction. Common examples include straight, curved, or turning path templates. Table 17 provides a summary of reviewed anchor-based prediction methods.
@@
Type Reference Description Goal Anchor Based Zhang et al. [527] (2020) Selects potential goal points based on observed trajectories and contextual features and ranks candidate goals based on learned scoring function. Tran et al. [528] (2020) Dual-channel framework: a goal channel predicts potential destinations, and a trajectory channel generates detailed paths conditioned on these goals. Mangalam et al. [529] (2020) Generates goals by sampling plausible future endpoints directly from the scene context, leveraging map features, and observed trajectories. Mangalam et al. [530] (2021) Hierarchically models goals, waypoints, and paths, generating scene-compliant multimodal predictions. Zhao and Wildes [531] (2021) Leverages Dynamic Time Warping (DTW) to retrieve pseudo goal positions from an expert repository of previously observed trajectories. Yao et al. [176] (2021) Uses a recognition network to encode the relationship between past trajectories and target goals, and a prior network for sampling. Wang et al. [532] (2022) Generates coarse stepwise goals iteratively, incorporates them using an attention-based aggregator. Chiara et al. [282] (2022) U-Net-based architecture to predict future goals as probability distributions based on scene semantics. Wang et al. [533] (2023) Uses a ”GoICrop” module to aggregate map features around potential goal areas, incorporating spatial interactions between agents and map. Dong et al. [534] (2023) Learns goal candidates by identifying representative points from future trajectories, using a learned mask conditioned on observed trajectories to guide selection. Bae and Jeon [535] (2023) Predicts a sequence of control points to divide pedestrian’s trajectory into three sections, capturing intermediate decisions. Wang et al. [525] (2023) Generates goal anchors via self-attention mechanism on the unified agent-centric scene representation, embedding contextual information. Chen et al. [518] (2024) Goals are predicted for each agent using a goal-guided state refinement (SSR) module. Trajectory Anchor Based Song et al. [511] (2021) Samples dynamic reference paths and imposes explicit kinematic and environmental constraints to produce feasibility-guaranteed trajectories. Varadarajan et al. [3] (2022) Learns context-aware latent anchor embeddings that adapt dynamically to the scene, guiding diverse and accurate trajectory predictions. Zhou et al. [526] (2023) Generates adaptive trajectory anchors via a proposal module, refined with context-specific biases through an anchor-based decoder.